Cornerstone Exercise 11.2 (Algorithmic)
Activity-Based Supplier Costing

Ventana Company is a car window repair and replacement company operating in the after-sales market. Ventana’s purchasing manager uses two suppliers (Jones Glass and Claro Glass) for the source of its passenger car windows. Data relating to side windows (Side) and windshields (WS) are given below.

I. Activity costs

Activity
Adverse buying* 697,500
Supplier returns** 74,100

* Extra cost of purchasing from local car dealer because of insufficient delivery of supplier.
** Windows returned because they were not ordered or because they were defective.

II. Supplier Data

Jones Glass Claro Glass
Side WS Side WS
Unit purchase price $70 $145 $67 $142
Units purchased 20,000 20,000 30,000 30,000
Insufficient units 1,050 1,050 3,600 3,600
Returned units 425 425 1,000 1,000

Required:

1. Calculate the activity rates for assigning costs to suppliers.

Adverse buying rate $ per adverse purchase
Supplier return rate $ per return

2. Calculate the total unit purchasing cost for each component for each supplier. If required, round your answers to the nearest cent.

Total unit cost
Jones Glass Claro Glass
Side $ $
WS $ $
Answer #1
Required 1 :
Total insufficient units = 1050 + 1050 + 3600 + 3600 9300
Total returned units = 425 + 425 + 1000 + 1000 2850
Adverse buying rate = Adverse buying cost / Total insufficient units = 697500 / 9300 75 per adverse purchase
Supplier return rate = Supplier return cost / Total returned units = 74100 / 2850 26 per return
Required 2 :
Allocation of activity cost :
Jones Glass Claro Glass
Side WS Side WS

Adverse buying cost

78750

[ 1050*75 ]

78750

[ 1050*75 ]

270000

[ 3600*75 ]

270000

[ 3600*75 ]

Supplier returns cost

11050

[ 425*26 ]

11050

[ 425*26 ]

26000

[ 1000*26 ]

26000

[ 1000*26 ]

Total activity cost 89800 89800 296000 296000
(/) Units purchased 20000 20000 30000 30000
Activity cost per unit 4.49 4.49 9.87 9.87

Total unit cost

[ Unit purchase price + Activity cost per unit ]

Jones Glass Claro Glass

Side

74.49

[ 70+4.49 ]

76.87

[ 67+9.87 ]

WS

149.49

[ 145+4.49 ]

151.87

[ 142+9.87 ]